Close Problens Calculate the pH of a weak acid solution (IHAb 100.Ka). Calculate the pH of a 0.406 M aqueous solution of nitrous acid (HNO,, K-4.5x10) and the equilibrium concentrations of the weak acid and its conjugate base. pHH [HNO2 lequilibrium NO2 lequilbrium Show Approach

media%2Fcec%2Fcec5a196-632e-4c59-9e26-31Help plz
Answer #1

Given

[HNO2] = 0.406 M

To calculate pH of a weak acid we have to use ICE table

HNO2 H2O <===> H3O+ NO2-
I 0.406 M - 0 0
C -x - +x +x
E 0.406-x - x x

Ka of HNO2 = 4.5 x 10-4

but

K_a =\frac{[H_3O^+][NO_2^-]}{[HNO_2]}

4.5\times 10^{-4} =\frac{[x][x]}{[0.406-x]}

4.5\times 10^{-4} =\frac{[x]^2}{[0.406-x]}

Neglecting denominator x (0.406-x = 0.406)

4.5\times 10^{-4} =\frac{[x]^2}{[0.406]}

4.5\times 10^{-4}\times 0.406 =[x]^2

1.827\times 10^{-4} =[x]^2

x=\sqrt{1.827\times 10^{-4}}

x = 0.0135 M

[H3O+] = x = 0.0135 M

[NO2-] = x= 0.0135 M

[HNO2] = 0406 - x = 0.406- 0.0135 = 0.392 M

Formula

pH = -log[H3O+]

pH = -log (0.0135)

pH = 1.87

------------------------------------------

Given,

[HC9H7O4] = 0.0170 M

Using ICE table to solve for pH

HC9H7O4 H2O <==> C9H7O4- H3O+
I 0.0170 M - 0 0
C -x - +x +x
E 0.0170 -x - x x

Given

K_a = \frac{[C_9H_7O_4^-][H^+]}{[HC_9H_7O_4]}

3.4\times 10^{-4}= \frac{[x][x]}{[0.0170-x]}

3.4\times 10^{-4}\times [0.0170-x]= [x]^2

here we are asked to solve by quadratic equation so we cannot neglect x

x^2=5.78\times 10^{-6}-3.4\times 10^{-4}(x)

x^2+3.4\times 10^{-4}(x)+(-5.78\times 10^{-6})=0

it is in the form of ax2 +bx +c =0

formula to solve for x

x = \frac{-b\pm \sqrt{b^2-4ac}}{2a}

x = \frac{(-3.4\times 10^{-4})\pm \sqrt{(3.4\times 10^{-4})^2-4\times (-5.78\times 10^{-6})}}{2}x = \frac{(-3.4\times 10^{-4})\pm \sqrt{2.32\times 10^{-5}}}{2}

x = \frac{(-3.4\times 10^{-4})\pm 0.00482}{2}

x = \frac{(-3.4\times 10^{-4})+ 0.00482}{2}, x = \frac{(-3.4\times 10^{-4})- 0.00482}{2}

x = 0.00224, x = -0.00258

Concentration wont be in negative

so X= 0.00224 M

[H3O+] = 0.00224 M

[C9H7O4-] = x =  0.00224 M

[HC9H7O4] = 0.0170- x = 0.0170- 0.00224 = 0.0148 M

pH = -log[H3O+]

pH = -log(0.00224)

pH = 2.65